My little brother has this on his HW and both of us are stuck:
Anne has $2.00 in nickels and dimes. She has 5 more dimes than nickels. How many dimes and nickels does she have?

I kept getting 0=0 and then tried this [d=dimes,n=nickels]: {2.00=.10d+.05n} {d=n+5}

Answer by scott8148(6628)

you have the right equations, just solve the system

200 = 10d + 5n

d = n + 5 ___ 10d = 10n + 50

200 = (10n + 50) + 5n ___ 150 = 15n